# Yogesh Kulkarni, Pooyan Fazli

### v1

@@ Acknowledgments
Appendix A Appendix
GRPO for Multimodal Reasoning
Recent works increasingly use RL methods, particularly GRPO (Xiao et al. 2025; Wan et al. 2025; Wang et al. 2025d), introducing new challenges related to training efficiency and stability. Wei et al. propose a “cold start” strategy via Supervised Fine-Tuning (SFT) before RL (Wei et al. 2025), while other approaches improve data efficiency through direct RL with test-time scaling  (Wang et al. 2025e). Concurrent advances enhance GRPO through curriculum learning (Lu et al. 2025), variance-aware data selection (Li et al. 2025a), hint-based guidance for difficult samples (Huang et al. 2025), and combining GRPO and DPO (Li et al. 2025d). Key limitations remain, including the “overthinking” issue, addressed by regulating reasoning length (Wang et al. 2025c; Xiao et al. 2025), and perception errors mitigated via grounded reasoning approaches (Fan et al. 2025b; Sarch et al. 2025) and new RL objectives such as PAPO’s Implicit Perception Loss (Wang et al. 2025d).
Audio-visual reasoning that incorporates audio often involves trade-offs between temporal coverage and spatial resolution. These challenges have been addressed through two-system architectures (Zhong et al. 2025), explicit contextual rewards (Zhao et al. 2025a), and task-specific improvements (Lu et al. 2025; Xing et al. 2025). AVATAR introduces a unified framework that integrates a stratified replay buffer for curriculum learning with Temporal Advantage Shaping (TAS) for credit assignment. This approach improves reasoning and grounding within a single model, eliminating the need for two-system architectures (Zhong et al. 2025) or explicit visual coordinate generation (Sarch et al. 2025).
@@ Appendix A Appendix
Figure 5: Performance vs. on-policy/off-policy sample ratio across six benchmarks. AVATAR achieves optimal performance with 4-4 split (4 on-policy, 4 off-policy samples), demonstrating that balanced mixing prevents both policy drift from excessive off-policy data and sample inefficiency from pure on-policy training.
Ablation Configuration Audio-Visual Benchmarks Video Reasoning Benchmarks
Group Setting OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO
Baseline Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0
Reward Suite $R_{\text{acc}}+R_{\text{format}}$ 46.8+2.6 45.2+1.2 30.1+0.3 42.3+1.7 62.1+1.9 28.9-0.1
$R_{\text{acc}}+R_{\text{format}}+R_{\text{self}}$ 47.5+3.3 45.8+1.8 30.7+0.9 43.1+2.5 63.4+3.2 29.4+0.4
$R_{\text{acc}}+R_{\text{format}}+R_{\text{self}}+R_{\text{judge}}$ 48.3+4.1 46.4+2.4 31.4+1.6 44.2+3.6 64.7+4.5 30.1+1.1
$R_{\text{acc}}+R_{\text{format}}+R_{\text{self}}+R_{\text{judge}}+VCRS$ 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
@@
Table 6: Ablation studies on our reward strategy across three training stages. Stage 1 uses $R_{\text{acc}}$ and $R_{\text{format}}$ for basic accuracy and format compliance. Stage 2 adds $R_{\text{self}}$ (majority vote consensus) to the existing rewards, enabling consistency learning. Stage 3 incorporates $R_{\text{judge}}$ (stepwise VLM evaluation) alongside all previous rewards, providing detailed reasoning feedback. VCRS maintains stable advantage baselines through moving averages, preventing vanishing advantages. Each cumulative addition demonstrates clear performance gains, with largest improvements on reasoning-heavy benchmarks (Video-Holmes: +4.5, MMVU: +5.4) where structured feedback proves most beneficial.
Ablation Studies
What is the impact of the on-policy/off-policy sample ratio?
@@
To determine the optimal balance between on-policy and off-policy data, we analyze performance across different sample ratios within a fixed group size of eight (Figure 5). We observe a consistent trend across all benchmarks: performance peaks when using a balanced 4-4 split (4 on-policy, 4 off-policy samples). Using fewer off-policy samples (0-3) fails to fully mitigate the data inefficiency of a purely on-policy approach. On the other hand, over-reliance on off-policy samples (5-7) degrades performance due to instability, as the policy drifts too far from the older, “stale” data in the replay buffer. The 4-4 split therefore represents an empirical sweet spot for AVATAR, maximizing the data efficiency gains of our off-policy architecture while maintaining the training stability necessary for effective learning.
@@
How do isolating rewards impact performance?
Table 6 shows the cumulative impact of each reward component in AVATAR’s training pipeline. The baseline, which includes only accuracy and format rewards, offers limited learning signals, especially on complex reasoning tasks such as Video-Holmes and MMVU, where binary feedback proves insufficient. Introducing self-rewarding ($R_{\text{self}}$) in Stage 2 enables consensus-based learning from the model’s own outputs, yielding moderate improvements on audio-visual tasks that benefit from internal consistency signals (e.g., +0.7 on OmniBench, +1.3 on MMVU). Adding the stepwise judge reward ($R_{\text{judge}}$) in Stage 3 results in larger gains by assessing intermediate reasoning steps, particularly enhancing performance on fine-grained localization benchmarks (e.g., +0.8 on OmniBench, +1.3 on MMVU). Finally, incorporating VCRS further stabilizes training by maintaining non-zero advantage baselines via moving averages, allowing the full reward setup to achieve the highest observed performance (e.g., +0.8 on OmniBench, +0.9 on MMVU).
How do training dynamics differ across benchmarks?
@@
In contrast, AVATAR shows a dip-and-recovery pattern that reflects the interplay of its key components. The initial drop in performance, especially on challenging benchmarks like OmniBench, Video-Holmes, and MMVU, occurs when the stratified replay buffer transitions from easy to hard samples, exposing the model to its most difficult failures. This performance dip would persist without TAS; however, the U-shaped temporal weighting in TAS provides focused learning signals that extract meaningful gradients from these difficult examples. By emphasizing the planning and synthesis phases that uniform credit assignment would dilute, TAS enables recovery and subsequent improvement. The model’s rebound and higher final performance confirm that our off-policy architecture transforms early failure into learning opportunities through repeated exposure, while TAS ensures each replayed experience yields maximum value. Notably, these gains emerge within just 1000 iterations, underscoring AVATAR’s superior sample efficiency and training effectiveness.
Does AVATAR generalize to smaller models?
@@
To evaluate the generalizability of our approach across different model scales, we conduct experiments with a 3B parameter version of the base Qwen2.5 Omni (Xu et al. 2025) model. Tables 7 and 8 demonstrate that AVATAR consistently outperforms both the baseline and on-policy GRPO across all benchmarks, achieving notable improvements of +3.4 on OmniBench, +2.9 on Video-Holmes, and +2.8 on MMVU. These results validate that AVATAR’s core components, the stratified replay buffer and TAS, effectively enhance multimodal reasoning capabilities regardless of language model scale. The consistent performance patterns across different model sizes further confirm that our off-policy architecture is broadly applicable, with particularly strong benefits observed on complex reasoning tasks where AVATAR’s targeted learning signals prove most impactful for guiding model improvement.
Audio-Visual Benchmarks
Model OmniBench DailyOmni AV-Counting AV-Odyssey WorldSense IntentBench
Baseline 3B Omni 42.4 42.9 19.7 27.2 35.2 59.7
Baseline GRPO 43.1+0.7 43.4+0.5 20.0+0.3 28.5+1.3 35.8+0.6 59.9+0.2
AVATAR 3B 45.8+3.4 44.7+1.8 20.9+1.2 29.8+2.6 37.1+1.9 60.5+0.8
@@
Table 7: Audio-visual benchmarks: 3B model comparison. AVATAR 3B consistently outperforms both the baseline 3B Omni and baseline GRPO across all benchmarks, achieving improvements of +3.4 on OmniBench and +2.6 on AV-Odyssey. The results demonstrate that our off-policy architecture and TAS effectively enhance multimodal reasoning capabilities even with smaller model scales, validating the broad applicability of our approach. Performance improvements over baseline 3B Omni are shown as subscripts.
General Video Understanding Video Reasoning
Model MVBench Video-MME LVBench Video-Holmes MMVU TOMATO
Baseline 3B Omni 59.2 57.6 34.3 39.2 62.0 27.0
Baseline GRPO 59.8+0.6 58.1+0.5 34.7+0.4 40.8+1.6 63.2+1.2 27.3+0.3
AVATAR 3B 61.2+2.0 59.9+2.3 35.8+1.5 42.1+2.9 64.8+2.8 28.4+1.4

### v2

@@ Acknowledgments
Appendix
Appendix A GRPO for Multimodal Reasoning
Recent works increasingly use RL methods, particularly GRPO [fastslow, srpo, papo], introducing new challenges related to training efficiency and stability. Wei et al. propose a “cold start” strategy via Supervised Fine-Tuning (SFT) before RL [wei2025advancing], while other approaches improve data efficiency through direct RL with test-time scaling  [videorts]. Concurrent advances enhance GRPO through curriculum learning [avreasoner], variance-aware data selection [li2025reinforcement], hint-based guidance for difficult samples [hintgrpo], and combining GRPO and DPO [veripo]. Key limitations remain, including the “overthinking” issue, addressed by regulating reasoning length [pixelthink, fastslow], and perception errors mitigated via grounded reasoning approaches [grit, grounded] and new RL objectives such as PAPO’s Implicit Perception Loss [papo].
Audio-visual reasoning that incorporates audio often involves trade-offs between temporal coverage and spatial resolution. These challenges have been addressed through two-system architectures [omnir1], explicit contextual rewards [humanomni], and task-specific improvements [avreasoner, echoink]. AVATAR introduces a unified framework that integrates a stratified replay buffer for curriculum learning with Temporal Advantage Shaping (TAS) for credit assignment. This approach improves reasoning and grounding within a single model, eliminating the need for two-system architectures [omnir1] or explicit visual coordinate generation [grounded].
Appendix B Ablation Studies
RQ1: Is TAS a principled, critic-free credit assignment strategy? We argue that TAS is a data-driven approximation that directly exploits the strong structural prior learned during our SFT (Stage 0). The SFT phase trains the model to generate reasoning in a predictable planning-evidence-synthesis structure.
To validate this, we empirically analyzed the positional distribution of “critical reasoning steps” (tokens with the highest causal impact on the final answer) from $50,000$ samples in our SFT dataset. This analysis was verified using GPT-4o [hurst2024gpt] (Prompt provided in Figure 12) to identify the most critical reasoning step for each sample. Figure 6 shows this empirical distribution (blue bars), which is distinctly parabolic-shaped: critical tokens cluster at the sequence start (planning, $\tilde{t}\in[0,0.2]$) and end (synthesis, $\tilde{t}\in[0.8,1.0]$), while intermediate evidence gathering tokens are far less critical.
Figure 6: TAS is a Data-Driven Approximation. We empirically analyze the distribution of “critical reasoning steps” from our SFT data. The resulting U-shaped histogram (blue bars) shows that important tokens cluster at the planning and synthesis phases. TAS’s parabolic function (orange, $\lambda_{\mathrm{TAS}}{=}0.3$) is a simple, critic-free function that directly approximates this observed data.
This empirical result validates that TAS’s parabolic function (orange line) is a principled, low-cost approximation of the true importance distribution in our data. This design is a key advantage: it achieves effective temporal credit assignment by leveraging the known structure of the reasoning, bypassing the need to train a complex, unstable, and high-variance value function or critic. Standard GRPO’s uniform credit assignment dilutes the learning signal by applying $A_{i}$ to all tokens equally. By modulating this with our data-driven weights $w_{i,t}$, the shaped advantage $A_{i,t}^{TAS}$ focuses the policy gradient onto the tokens most critical for a correct outcome, providing a stable and efficient learning signal.
@@ Appendix B Ablation Studies
Figure 7: Performance vs. on-policy/off-policy sample ratio across six benchmarks. AVATAR achieves optimal performance with 44-44 split (44 on-policy, 44 off-policy samples), demonstrating that balanced mixing prevents both policy drift from excessive off-policy data and sample inefficiency from pure on-policy training.
Figure 8: TAS Empirically Focuses the Learning Signal. We plot the average per-token loss (the learning signal) vs. normalized position across the Video-Holmes [videoholmes] benchmark. The Baseline GRPO (blue dashed line) shows a flat, unfocused signal, indicative of uniform credit assignment. AVATAR w/ TAS (orange solid line) successfully focuses the learning signal, amplifying it on the critical planning and synthesis phases.
RQ2: Does TAS empirically focus the learning signal in practice?
Figure 6 established why a parabolic-shaped function is the correct model: the problem (critical steps) is empirically parabolic-shaped. Here, we validate that our solution (the learning signal) successfully adopts this parabolic-shape in practice. We performed an analysis over the Video-Holmes [videoholmes] benchmark, calculating the average per-token loss (the learning signal) against its normalized position. Figure 8 shows the result. The Baseline GRPO (blue dashed line) exhibits a relatively flat loss distribution. This is the empirical evidence of uniform credit assignment: the learning signal is diluted by being applied unfocused and equally to all tokens, regardless of their importance. In contrast, AVATAR w/ TAS (orange solid line) demonstrates a clear parabolic-shaped loss profile. The shaped advantage $A_{i,t}^{TAS}$amplifies the learning signal at the critical planning ($\tilde{t}\in[0,0.2]$) and synthesis ($\tilde{t}\in[0.8,1.0]$) phases. This proves TAS is a functioning mechanism: it forces the optimizer to learn most aggressively from errors in the most important parts of the reasoning chain.
Config Halluc TP AC Baseline GRPO 35.6 53.6 32.5 AVATAR w/o Off-Policy 40.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5})}$ 54.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ 34.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.6})}$ AVATAR w/o TAS 48.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+13.3})}$ 56.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.5})}$ 35.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.3})}$ AVATAR (Full) 51.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+15.6})}$ 57.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.6})}$ 37.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.8})}$
@@
Table 8: Ablation studies on our reward strategy across three training stages. Stage 11 uses $R_{\mathrm{acc}}$ and $R_{\mathrm{format}}$ for basic accuracy and format compliance. Stage 22 adds $R_{\mathrm{self}}$ (majority vote consensus) to the existing rewards, enabling consistency learning. Stage 33 incorporates $R_{\mathrm{judge}}$ (stepwise VLM evaluation) alongside all previous rewards, providing detailed reasoning feedback. VCRS maintains stable advantage baselines through moving averages, preventing vanishing advantages. Each cumulative addition demonstrates clear performance gains, with largest improvements on reasoning-heavy benchmarks (Video-Holmes: +4.5+4.5, MMVU: +5.4+5.4) where structured feedback proves most beneficial.
Ablation Configuration Audio-Visual Benchmarks Video Reasoning Benchmarks Group Setting OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO Baseline Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0 Reward Suite $R_{\mathrm{acc}}+R_{\mathrm{format}}$ 46.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.6})}$ 45.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ 30.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.3})}$ 42.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.7})}$ 62.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.9})}$ 28.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{-0.1})}$ $R_{\mathrm{acc}}+R_{\mathrm{format}}+R_{\mathrm{self}}$ 47.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.3})}$ 45.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ 30.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.9})}$ 43.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.5})}$ 63.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.2})}$ 29.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.4})}$ $R_{\mathrm{acc}}+R_{\mathrm{format}}+R_{\mathrm{self}}+R_{\mathrm{judge}}$ 48.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.1})}$ 46.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.4})}$ 31.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.6})}$ 44.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.6})}$ 64.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5})}$ 30.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.1})}$ $R_{\mathrm{acc}}+R_{\mathrm{format}}+R_{\mathrm{self}}+R_{\mathrm{judge}}+\mathrm{VCRS}$ 49.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.9})}$ 47.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.0})}$ 32.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3})}$ 45.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5})}$ 65.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+5.4})}$ 30.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$
RQ3: What is the impact of the on-policy/off-policy sample ratio?
@@
To determine the optimal balance between on-policy and off-policy data, we analyze performance across different sample ratios within a fixed group size of eight (Figure 7). We observe a consistent trend across all benchmarks: performance peaks when using a balanced 44-44 split (44 on-policy, 44 off-policy samples). Using fewer off-policy samples (0-33) fails to fully mitigate the data inefficiency of a purely on-policy approach. On the other hand, over-reliance on off-policy samples (55-77) degrades performance due to instability, as the policy drifts too far from the older, “stale” data in the replay buffer. The 44-44 split therefore represents an empirical sweet spot for AVATAR, maximizing the data efficiency gains of our off-policy architecture while maintaining the training stability necessary for effective learning.
RQ4: How do AVATAR’s components impact audio-visual reasoning?
RQ5: Impact of $\lambda_{\mathrm{TAS}}$ on Learning Stability.
As illustrated in Figure 9, model performance exhibits a clear three-phase trend with respect to $\lambda_{\mathrm{TAS}}$. At low values ($\lambda_{\mathrm{TAS}}{<}0.1$), the reward shaping effect is weak, leading to slower convergence and unstable gains. As $\lambda_{\mathrm{TAS}}$ increases to a moderate level ($\lambda_{\mathrm{TAS}}{\approx}0.3$), performance peaks across both audio-visual and video reasoning benchmarks, reflecting a balanced trade-off between stability and exploration. However, at higher values ($\lambda_{\mathrm{TAS}}{>}1.0$), the model becomes over-regularized, reducing adaptability and leading to gradual accuracy decline.
RQ6: How do isolating rewards impact performance?
Table 8 shows the cumulative impact of each reward component in AVATAR’s training pipeline. The baseline, which includes only accuracy and format rewards, offers limited learning signals, especially on complex reasoning tasks such as Video-Holmes and MMVU, where binary feedback proves insufficient. Introducing self-rewarding ($R_{\mathrm{self}}$) in Stage 22 enables consensus-based learning from the model’s own outputs, yielding moderate improvements on audio-visual tasks that benefit from internal consistency signals (e.g., +0.7+0.7 on OmniBench, +1.3+1.3 on MMVU). Adding the stepwise judge reward ($R_{\mathrm{judge}}$) in Stage 33 results in larger gains by assessing intermediate reasoning steps, particularly enhancing performance on fine-grained localization benchmarks (e.g., +0.8+0.8 on OmniBench, +1.3+1.3 on MMVU). Finally, incorporating VCRS further stabilizes training by maintaining non-zero advantage baselines via moving averages, allowing the full reward setup to achieve the highest observed performance (e.g., +0.8+0.8 on OmniBench, +0.9+0.9 on MMVU).
RQ7: How do training dynamics differ across benchmarks?
(a) Audio-Visual Reasoning
(b) Video Reasoning
Figure 9: Performance variation with TAS weighting $(\lambda_{\mathrm{TAS}})$. Both (a) Audio-Visual and (b) Video Reasoning benchmarks peak around $\lambda_{\mathrm{TAS}}{=}0.3$, showing improved stability and reward shaping before slight decline at higher values.
@@
(a) OmniBench
(b) DailyOmni
(c) AV-Odyssey
(d) WorldSense
(e) IntentBench
(f) Video-Holmes
(g) MMVU
(h) TOMATO
(i) LVBench
Figure 11: Ablation on total replay buffer capacity across audio-visual and video reasoning benchmarks. Performance peaks consistently at around 10k buffer size, balancing diversity and staleness in experience replay. AVATAR achieves higher and more stable gains than GRPO across all tasks, validating its curriculum-based replay and TAS advantage shaping as critical for robust multimodal learning.
@@
In contrast, AVATAR shows a dip-and-recovery pattern that reflects the interplay of its key components. The initial drop in performance, especially on challenging benchmarks like OmniBench, Video-Holmes, and MMVU, occurs when the stratified replay buffer transitions from easy to hard samples, exposing the model to its most difficult failures. This performance dip would persist without TAS; however, the U-shaped temporal weighting in TAS provides focused learning signals that extract meaningful gradients from these difficult examples. By emphasizing the planning and synthesis phases that uniform credit assignment would dilute, TAS enables recovery and subsequent improvement. The model’s rebound and higher final performance confirm that our off-policy architecture transforms early failure into learning opportunities through repeated exposure, while TAS ensures each replayed experience yields maximum value. Notably, these gains emerge within just 10001000 iterations, underscoring AVATAR’s superior sample efficiency and training effectiveness.
RQ8: What is impact of replay buffer size?
The total size of the replay buffer introduces a critical trade-off between sample diversity and data staleness. A larger buffer holds a more diverse set of past experiences, especially difficult samples, which can improve generalization. However, a buffer that is too large increases the risk of sampling “stale” data generated by a much older policy, which can destabilize the off-policy learning. This trade-off is empirically validated in Figure 11. Across all benchmarks, performance consistently peaks at an optimal capacity of 10k samples. Below this size, performance suffers from a lack of diversity; beyond this point, performance degrades as the negative impact of data staleness begins to outweigh the benefits of diversity.
RQ9: Does AVATAR generalize to smaller models?
@@
To evaluate the generalizability of our approach across different model scales, we conduct experiments with a $3B$ parameter version of the base Qwen2.5 Omni [qwen25omni] model. Tables 9 and 10 demonstrate that AVATAR consistently outperforms both the baseline and on-policy GRPO across all benchmarks, achieving notable improvements of +3.4+3.4 on OmniBench, +2.9+2.9 on Video-Holmes, and +2.8+2.8 on MMVU. These results validate that AVATAR’s core components, the stratified replay buffer and TAS, effectively enhance multimodal reasoning capabilities regardless of language model scale. The consistent performance patterns across different model sizes further confirm that our off-policy architecture is broadly applicable, with particularly strong benefits observed on complex reasoning tasks where AVATAR’s targeted learning signals prove most impactful for guiding model improvement.
@@
Table 9: Audio-visual benchmarks: $3B$ model comparison. AVATAR 3B consistently outperforms both the baseline $3B$ Omni and baseline GRPO across all benchmarks, achieving improvements of +3.4+3.4 on OmniBench and +2.6+2.6 on AV-Odyssey. The results demonstrate that our off-policy architecture and TAS effectively enhance multimodal reasoning capabilities even with smaller model scales, validating the broad applicability of our approach. Performance improvements over baseline 3B Omni are shown in green.
Audio-Visual Benchmarks Model OmniBench DailyOmni AV-Counting AV-Odyssey WorldSense IntentBench Baseline $3B$ Omni 42.4 42.9 19.7 27.2 35.2 59.7 Baseline GRPO 43.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.7})}$ 43.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.5})}$ 20.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.3})}$ 28.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.3})}$ 35.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.6})}$ 59.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.2})}$ AVATAR $3B$ 45.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.4})}$ 44.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ 20.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ 29.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.6})}$ 37.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.9})}$ 60.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.8})}$